# PEEB: Part-based Image Classifiers with an Explainable and Editable Language Bottleneck

### v1

@@ 4.3 Training strategy
We empirically find that solely training the part MLP layer does not achieve the desired classification accuracy, prompting us to update the image encoder. However, re-training this encoder impacts the linear projection and box MLP layers. As a result, we have to train all components together. All the components are initialized from the corresponding components in OWL-ViT, except for our proposed part MLP. Our training strategy has two phases: two-stage pre-training on the large-scale Bird-11K dataset and fine-tuning on downstream tasks. We provide full hyperparameter details of PEEB in Sec. A.8.
@@
Objectives There are three objectives to train PEEB: (a) Train the part MLP layer contrastively to maximize the similarity between related part-descriptor pairs while minimizing the unrelated pairs using symmetric cross-entropy (CE) loss (Radford et al., 2021); (b) Train the linear projection layer to mimic OWL-ViT’s behaviors (i.e. the similarity matrix) for part selection with symmetric CE loss; and (c) Train the box MLP layer for bounding box regression with DETR losses (Zheng et al., 2020) i.e. a linear combination of corner-to-corner distance loss and GIoU loss (Rezatofighi et al., 2019). We provide detailed loss functions in Sec. A.10.
Challenges One problem emerges when we jointly train all components together: the model learns at a significantly slow pace since PEEB needs to learn to optimize two symmetric cross-entropy losses while maintaining the high-quality predicted boxes. To tackle this problem, we split the pre-training phase into two stages: (1) train the image encoder and part MLP layer with the first objective; then (2) train the linear projection and box MLP layers with the second and third objectives to accordingly adjust their weights to the changes in the image encoder. Notably, the text encoder is always frozen as it was designed for open vocabulary, its generalizability to unseen texts (i.e., descriptors of an unseen bird) should be preserved.
4.3.1 Pre-training on Bird-11K dataset
@@
Stage 1, contrastive learning with teacher model: The image encoder and part MLP layer are jointly trained using the symmetric CE loss, which is particularly suitable for PEEB as it learns the mapping between visual parts and descriptors. In this stage, we follow the teacher model – OWL-ViT to select part embeddings to ensure the selected part embeddings are semantically meaningful, e.g., can be used for box prediction (Fig. A2).
Stage 2, eliminate teacher dependency:
As the image encoder is modified in stage 1, the linear project and the box MLP needs to be updated accordingly. Specifically, we consider the teacher model OWL-ViT as ground-truths and train the linear projection layer using symmetric CE loss (Fig. A3, 1a–c, 2a–c). For box MLP, given the absence of human-annotated boxes for individual parts, we obtain pseudo labels sourced from OWL-ViT as ground-truths for the training with DETR losses (Fig. A3, 2d). In this training step, the image encoder is frozen while the part MLP layer is not involved. After two-step training, PEEB can perform zero-shot classification while providing the mappings between visual parts and descriptors as faithful explanations.
4.3.2 Fine-tuning on target datasets
@@
We can further fine-tune the pre-trained model on downstream tasks, e.g., CUB, NABirds and iNaturalist to compare with other baselines (e.g. prototype-based approaches). In this phase, to adapt to the downstream tasks, all components except the text encoder are trained jointly and the loss function for part MLP is changed from symmetric CE to CE while other losses are kept intact.
@@ 5 Experiments & Results
We conduct systematic experiments to compare the generalization ability of PEEB with explainable methods on two zero-shot settings: GZSL (Sec. 5.1 – 5.3) and ZSL (Sec. 5.4). Notably, our method demonstrates significantly superior performance in GZSL and generalizes better in ZSL setting.
Moreover, following Donnelly et al., 2022, we finetune and evaluate PEEB on downstream tasks to measure the transferability compared to other part- and text concept-based methods (Sec. 5.5). Results show that PEEB achieves decent performance while providing more faithful explanations (Figs. 4, A10, A11 and A12). According to the evaluation for visual part localization (Sec. E.7) and qualitative analysis (Appendix G), we find that the box prediction performance of PEEB is comparable with OWL-ViT.
Original Descriptors (a)Incorrect Descriptors (b)
@@
5.1 CLIP-based classifiers depend mostly on class names (not part descriptors)
M&V shows that incorporating descriptors generated by GPT-3 (Brown et al., 2020) to class names increases CLIP accuracy on downstream tasks. Yet, it remains unknown how useful the descriptors are compared to the randomized descriptors for M&V’s method.
Experiment To address the concern, we conduct two systematic studies: (1) randomly swap a set of part descriptors among classes in CUB, NABirds, and iNaturalist datasets and measure the contribution of descriptors by the difference in model performance on CUB test set (i.e., 200-way classification) using original and random descriptions. (2) Replace the common names with scientific names in CUB, NABirds, and iNaturalist and measure the difference in model performance. The first experiment aims to study the contributions of descriptions to the CLIP-based method, and the second experiment tries to understand the dependencies of the class names.
Results When random descriptors are used, M&V ’s accuracy drops marginally by -0.9 points, while PEEB’s performance significantly deteriorates, indicating a strong dependence on accurate descriptions. However, when only descriptors are used without class names (Table 1) or class names are replaced by scientific names (Table 2), M&V ’s accuracy reduces drastically to nearly random chance, underscoring the reliance of CLIP-based classifiers on class names.
5.2 Part-based pre-trained PEEB outperforms CLIP-based classifiers
@@
The dependence on class names suggests CLIP is potentially exposed to class names during training. Thus, we compare PEEB with the CLIP-based classifiers in the GZSL setting for a fair comparison.
Experiments We conduct two-stage pre-training for PEEB on Bird-11K where CUB, NABirds, and iNaturalist test sets are excluded for evaluation. Note that the contrastive pre-training is on part level, that is, PEEB does not have direct access to the class labels.
@@
Results PEEB outperforms the baselines across all three datasets, achieving improvements of (+10 to +12 points), (+28 to +29 points) and (+8 to +9 points) on CUB, NABirds and iNaturalist, respectively (Table 2).
| CLIP | M&V | PEEB (ours) | ||
| Using class names | ✗ | ✗ | ||
| Original Descriptors | 52.02 | 53.78 | 5.89 | 64.33 |
| Incorrect Descriptors | n/a | 52.88 | 0.59 | 0.88 |
5.3 PEEB is superior to text concept-based classifiers on GZSL setting
The advancements in Large Language Models (LLMs) led to the emergence of text concept-based approaches (Yuksekgonul et al., 2023; Yan et al., 2023; Esfandiarpoor and Bach, 2023; Panousis et al., 2023) aiming to provide a more flexible and explainable classifier. We compare PEEB in the GZSL setup on CUB test set with text concept-based methods.
Experiment We conduct an experiment with the same setting as in Sec. 5.2. While PEEB does not use specific concepts, the descriptors it utilizes (described in Appendix C) function as open-vocabulary concepts. Most text concept-based methods in GZSL require pre-defined concept sets, limiting their applicability to similar datasets like NABirds or iNaturalist. Hence, our evaluation is focused solely on the CUB dataset.
Results PEEB exhibits superior GZSL performance, outperforming recent text concept-based approaches by +3 to +10 points (Table 3). Unlike the traditional approaches, which primarily rely on predefined concept pools, PEEB utilizes domain-specific, natural language descriptors. This distinctive approach enables PEEB to adapt more flexibly to various datasets such as NABirds and iNaturalist. In contrast, concept pool-based methods often necessitate manual concept updates and potential re-training for such extensions.
| Method | CUB | Concept-type |
| FuDD (2023) | 54.30 | Class-wise Captions |
| Han et al. (2023) | 56.13 | Evolving Descriptions |
| PCBM (2023) | 61.00 | Pre-defined Concept Pool |
| Yan et al. (2023) | 60.27* | Pre-defined Concept Pool |
| LaBo (2023) | 54.19 | Pre-defined Concept Pool |
| PEEB (Ours) | 64.33 | Class-wise Descriptors |
@@ 5.4 PEEB generalizes to traditional ZSL
Some vision-language approaches enable the model to process certain textual-based concepts Ji et al. (2018); Zhu et al. (2018); Paz-Argaman et al. (2020); Chen et al. (2020); Kousha and Brubaker (2021); Rao et al. (2023), and therefore, ZSL is used to measure their generalization capability. In this evaluation, we employed the traditional ZSL split on CUB from Akata et al. and the Super-Category-Similar/Exclusive (SCS/SCE) proposed by Elhoseiny et al. on CUB and NABirds. The SCS (easy) and SCE (hard) split are designed based on the species hierarchy to intentionally have two levels of difficulties from the ZSL test.
Experiment Adhere to the conventional ZSL setting; we exclude all the CUB or NABirds classes for pre-training and fine-tune the model following Akata et al. or Elhoseiny et al.’s split. We randomly select 10% from the training set as the validation set and choose the checkpoints based on the lowest validation loss.
Results PEEB outperforms all baselines across 3 test splits (from CUB and NABirds) by (+4 to +10 points) in terms of harmonic mean, indicating that PEEB is more generalized to not only seen classes (80.78 vs 65.80) but also unseen classes (all other results in Table 4). The easy tests (SCS) guarantee the presence of classes similar to (but distinct from) the ones in the training set. Therefore, all baselines that learn better on the training set tend to have better accuracy in the easy test. Conversely, the hard splits (SCE) ensure the test classes are from different categories of the training classes. This distinction makes the hard test a more accurate metric for assessing a model’s generalization ability. PEEB excels over all baselines by (+5 to +15 points) (accuracy) for SCE split and +2.64 points (accuracy) compared to CLORE.
Furthermore, we also evaluate the model that is pre-trained without all CUB and NABirds classes and compare them with CLIP and M&V methods on CUB and NABirds. Interestingly, PEEB outperforms both baselines with (+10 to +12) points on CUB and (+1 to +3) points on NABirds (Sec. E.1). This finding further substantiates the generalization capability of our method.
@@
| Seen | Unseen | Harmonic | Seen | Unseen | Harmonic | |
| Data split by Akata et al. (2015) | ||||||
| CLORE (2022) | 65.80 | 39.10 | 49.05 | n/a | ||
@@
| SCS/SCE splits by Elhoseiny et al. (2017) | ||||||
| SCS | SCE | Harmonic | SCS | SCE | Harmonic | |
@@
5.5 Finetuning pre-trained PEEB on CUB-200 yields a competitive explainable classifier in supervised learning
We compare our model with explainable methods, such as ProtoPNet Chen et al. (2019), using the conventional pre-training and fine-tuning approach on Bird-11K and the CUB dataset. We focus on evaluating how effectively the pre-trained PEEB transfers to downstream tasks, providing insights into its adaptability and performance.
Experiment We further fine-tune all components of the pre-trained PEEB on the CUB dataset for comparison with other explainable methods. We fine-tune the pre-trained model for 30 epochs and select the best checkpoints based on validation loss. We provide the detailed hyperparameters in Table A2.
Results In the CUB dataset, PEEB trained from OWL-ViT backbone achieves 77.80% accuracy. With pre-training on Bird-11K, our model sets a new standard: 86.73% and 88.80% accuracy with two different backbones (Table 5), exceeding Deformable ProtoPNet (86.4%) and even ProtoTree (87.20%). PEEB distinctly maps descriptors to visual inputs, facilitating easier debugging and clearer understanding as demonstrated in Fig. 4.
| Methods | Model size | Backbone | Accuracy |
@@
| – Concept bottleneck classifiers | |||
| Concept Bottleneck Models (2020) | 11M | ResNet-18 (2016) | 80.10 |
| CPM (2023) | 155M | ViT-B/16 (2021) | 72.00 |
| CDM (2023) | 155M | ViT-B/16 | 74.31 |
| LaBo (2023) | 427M | ViT-L/14 | 81.90 |
| – Part-based classifiers | |||
@@
| ProtoTree (2021) | 92M* | ResNet-50 (2016) | 87.20* |
| TesNet (2021) | 79M | Dense121 (2017) | 84.80 |
@@
| PEEB (ours) – w/o pre-training | 155M | OWL-ViT | 77.80 |
| PEEB (ours) | 155M | OWL-ViT | 86.73 |
| PEEB (ours) | 155M | OWL-ViT | 88.80 |
5.6 Applying PEEB to Dog classification
PEEB can be applied across various domains, including dogs, cats, fish, and butterflies. We specifically demonstrate its applicability in the dog domain in this section.
Data preparation
We first define parts of interest that humans typically observe to distinguish dog breeds. Following PartImageNet He et al. (2021), we have 4 parts of interest: head, body, legs and tail and add two more parts: muzzle and ears as they also play a key role in dog recognition. Then, we prompt GPT-4 to get the descriptors for 6 parts per dog breed for 140 classes from Dog-140 – a large-scale dataset we construct from ImageNet-21K for pre-training (Sec. D.2). Finally, we run OWL-ViT model to detect the corresponding boxes for 6 pre-defined parts.
Experiment
Following bird experiments, we pre-train PEEB on Dog-140 and then further fine-tune on Stanford Dogs for evaluation. The experimental setup allows us to evaluate the pre-trained PEEB in GZSL setting and the fine-tuned PEEB on the downstream task. Additionally, we fine-tune PEEB from OWL-ViT backbone to get the lower bound performance on Stanford Dogs.
Results
PEEB achieves 74.17% accuracy on Stanford Dogs without pre-training on a large-scale dataset, which is similar to PEEB performance on CUB with 77.80% (Table 5). Gathering more images from ImageNet-21K for pre-training enhances PEEB’s accuracy to 87.38% for the pre-trained PEEB and 92.20% for fine-tuned PEEB (Table 6), while providing fine-grained explanations (Fig. 5). Interestingly, PEEB not only outperforms XAI methods but also achieves performance comparable to SOTA black boxes (92.20% vs 97.00%).
| Methods | Model size | Backbone | Accuracy |
| – Explainable methods | |||

### v3

@@ 4.3 Training strategy
Trainable networks In preliminary experiments, we find training only Part MLP (while keeping all other networks frozen) to result in poor accuracy. Therefore, we train Part MLP from scratch and also finetune the image encoder, Linear Projection, and Box MLP. We finetune all OWL-ViT components from their original weights. In contrast, our proposed Part MLP starts from random weights. Our training has two phases: (a) 2-stage pre-training on the large-scale Bird-11K dataset and (b) finetuning on downstream tasks. More hyperparameter details are in Sec. A.8.
@@
Objectives We aim to train PEEB to classify images well while maintaining the ability to detect object parts. This translates into three training objectives: (1) Train the Part MLP contrastively using a symmetric cross-entropy (SCE) loss (Radford et al., 2021) to maximize the similarity between region-text pairs while minimizing the similarity for negative pairs; (2) Train the Linear Projection using a SCE loss to mimic OWL-ViT’s behaviors (i.e. the similarity matrix) for part selection; and (3) Train Box MLP to predict bounding boxes with DETR losses (Zheng et al., 2021) i.e. a linear combination of corner-to-corner distance loss and GIoU loss (Rezatofighi et al., 2019).
All three losses are described in Sec. A.10.
A challenge when jointly minimizing all three losses above is that PEEB’s validation loss improves significantly slowly perhaps because of some tension between the two SCE losses and the DETR detection loss. To overcome this challenge, we split the pre-training phase into two stages: (1) first, train the image encoder and Part MLP for classification using the SCE loss; then (2) train the Linear Projection and Box MLP using the 2nd and 3rd loss so they can adapt their weights to the updated image encoder. We always keep the text encoder frozen since we want to preserve its generalizability to the descriptors of unseen objects.
4.3.1 2-stage pre-training on Bird-11K
@@
Stage 1: Contrastive learning The image encoder and Part MLP are jointly trained using a SCE loss, which allows PEEB to learn to map the visual parts to corresponding text descriptors. In this stage, we use a pre-trained OWL-ViT to select 12 part embeddings per input image (i.e., teacher forcing) to ensure the selection of part embeddings is meaningful and consistent while the embeddings themselves are updating (see Fig. A2).
Stage 2: Learning to detect from a teacher After the image encoder is modified in Stage 1, we then train the Linear Projection and Box MLP jointly. We use the OWL-ViT as the teacher to train both components. Using SCE loss, we train the Linear Projection such that the similarity matrix between the part-names and visual parts matches those of the teacher (Fig. A3, 1a–c, 2a–c). Given the absence of human-annotated boxes for object parts, we train Box MLP to predict the same boxes as the predictions by OWL-ViT using DETR losses (Fig. A3, 2d). In this Stage 2, the image encoder is frozen while Part MLP is not involved.
After 2-stage training, PEEB can perform zero-shot classification and generate explanations.
4.3.2 Finetuning on classification tasks
@@
We finetune the pre-trained PEEB on downstream tasks (CUB, NABirds and iNaturalist for birds and Dogs-120 for dogs) to further improve its accuracy. In this phase, to adapt to a downstream task, all components except the text encoder are trained jointly and the loss for Part MLP is changed from SCE (contrastive) to CE (classification) while the other two losses (DETR) are kept intact.
@@ 5 Experiments & Results
We conduct extensive experiments to evaluate PEEB on multiple bird datasets (CUB, NABirds, iNaturalist) and on GZSL (Secs. 5.1 and 5.3), ZSL (Sec. 5.4) and also supervised learning settings.
We also find PEEB to perform well on
dog image classification on Dogs-120 (Sec. 5.6).
Original Descriptors (a)Randomized, wrong descriptors (b)
@@
5.1 CLIP-based classifiers rely mostly on {class names} (not descriptors)
M&V show that inserting extra GPT-3-generated descriptors into CLIP’s prompts increases its accuracy on many tasks Menon and Vondrick (2023). Yet, it is unknown how important these expressive descriptors are compared to the class names. To answer this question, we conduct two experiments on all three models: CLIP, M&V, and our PEEB.
Experiment 1 We evaluate the role of expressive descriptors to CLIP-based models and to PEEB by measuring the drop in CUB-200 accuracy of each model when the descriptors are randomized.
For M&V and PEEB, we randomize the descriptors by swapping each descriptor with another from an arbitrary class (examples in Fig. 4).
Experiment 2 We test the dependence of models on class names by measuring the accuracy drop when they are replaced by scientific names (e.g., Painted Bunting Passerina ciris) on CUB, NABirds, and iNaturalist.
Results When random descriptors are used, M&V accuracy drops marginally by -0.9 pp (Table 1; 53.70% 52.88%), showing that descriptors actually play a minimal role in model predictions. Instead, CLIP and M&V mostly rely on class names (e.g., 53.78% 7.66%; Table 2)—their accuracy drops drastically when class names are replaced by scientific names, which are less common.
In contrast, the expressive part descriptors play a major role in PEEB whose accuracy decreases significantly to near random-chance (64.33% 0.88%; Table 1) when the descriptors are randomized. Indeed, in PEEB, the textual descriptors serve as editable and interpretable model parameters that can be refined and extended by humans to account for new classes (Fig. 2).
5.2 Pre-trained PEEB outperforms CLIP-based classifiers in GZSL
@@
The dependence on class names (Sec. 5.1) suggests that CLIP was exposed to these names during training. Thus, for a fair comparison, we compare PEEB with CLIP-based classifiers in the GZSL setting.
Experiment We train PEEB on Bird-11K using the two-stage pre-training (described in Sec. 4.3.1), and then test it on CUB, NABirds, and iNaturalist without any finetuning. That is, PEEB’s contrastive pre-training is at the part level and therefore the model has not seen the species labels of images.
@@
Results PEEB outperforms both CLIP and M&V on all three datasets by huge margins of around +10, +28, and +8 pp on CUB-200, NABirds-555 and iNaturalist-1486, respectively (see Table 2).
5.3 PEEB is superior to text descriptor-based classifiers in GZSL on CUB-200
The advent of CLIP (2021) by OpenAI enabled a class of image classifiers that match the input image with pre-defined textual prompts that may include class names or descriptors of the classes. Yet, in contrast to PEEB, these descriptors often describe the entire image and are also matched (via dot product) with the entire image instead of image regions. Here, we compare PEEB with these methods in the GZSL setting on CUB-200.
Experiment We repeat the same experiments in Sec. 5.2. As these bird classifiers (listed in Table 3) were reported on CUB only (not NABirds or iNaturalist), our comparison is on CUB.
Results PEEB exhibits superior GZSL performance, outperforming recent text concept-based approaches by +3 to +10 pp (Table 3b). Compared to prior methods, PEEB is the only one to detect visual object parts and match them with text descriptors. Furthermore, attribute-based classifiers, e.g., Yuksekgonul et al. (2023) require re-training to adapt to new classes or datasets (e.g., NABirds or iNaturalist) in the same domain. In contrast, to apply PEEB to NABirds or a new class, no training is required—it is necessary to only edit its text descriptors (see Fig. 2). Interestingly, PEEB is 2nd-best model, only after GPT-4V (64.33% vs. 69.40%), which is given the same textual part descriptors as PEEB for all 200 CUB classes and asked to select a matching class for each image.
* 1-shot learning. † -means with .
| Method | Acc (%) | {c} | Textual descriptors |
| (a) Vision-language models with class names {c} in the prompt | |||
| CLIP (2021) | 52.02 | ✓ | Image-level |
| M&V (2023) | 53.78 | ✓ | Image-level |
| FuDD (2023) | 54.30 | ✓ | Image-level |
| Han et al. (2023b) | 56.13 | ✓ | Image-level |
| (b) Vision-language models with text bottlenecks and no class names {c} | |||
| LaBo (2023) | 54.19† | ✗ | Image-level |
| Yan et al. (2023) | 60.27* | ✗ | Image-level, attribute-based |
| PEEB (ours) | 64.33 | ✗ | Part-level |
| GPT-4V (2023) | 69.40 | ✓ | Part-level |
| (c) Concept-Bottleneck Models with attribute-based, non-textual bottlenecks | |||
| CBM (2020) | 62.90 | ✗ | Attribute-based, tabular data |
| PCBM (2023) | 61.00 | ✗ | Attribute-based, tabular data |
@@ 5.4 PEEB generalizes to traditional ZSL
Since PEEB outperforms modern vision-language models in GZSL (Sec. 5.3), we are motivated to further compare PEEB with SOTA approaches in the traditional ZSL setting (where the test classes are excluded from all prior training).
Experiment We evaluate PEEB on two common ZSL splits: (a) the CUB split Akata et al. (2015); and (b) the Super-Category-Similar/Exclusive (SCS/SCE) splits Elhoseiny et al. (2017) on CUB and NABirds. The SCS (Easy) and SCE (Hard) splits are designed to test two generalization levels (generalizing to close vs. distant unseen species).
Aligned with ZSL conventions, we exclude all species that exist in CUB or NABirds from the pre-training and then finetune PEEB using the train/test splits by Akata et al. and Elhoseiny et al.. We randomly take 10% of the training set as the validation set and choose the checkpoints based on the lowest validation loss.
@@
| Seen | Unseen | Mean | Seen | Unseen | Mean | |
| (a) Data split by Akata et al. (2015) | ||||||
| CLORECLIP (2023a) | 65.80 | 39.10 | 49.05 | n/a | ||
@@
| (b) SCS/SCE splits by Elhoseiny et al. (2017) | ||||||
| SCS | SCE | Mean | SCS | SCE | Mean | |
@@
Results By a large margin, PEEB outperforms CLORECLIP, a SOTA CUB method in the (2015) split, on both seen and unseen classes (Table 4a). On the (2017) splits, PEEB is the SOTA in the Hard set on both CUB and NABirds datasets (Table 4b). That is, PEEB is better in generalizing to distant, unseen classes. This may be because PEEB decomposes both the image and the text descriptors into part-level features, which can re-combine to match an arbitrary unseen class (as illustrated in Fig. 2).
Interestingly, on both CUB and NABirds, PEEB is competitive but not SOTA on the Easy sets (Table 4b; Easy)—those classes that are close to the training-set classes and thus considered easier to identify. Overall, considering the harmonic mean over both Easy and Hard accuracy scores, PEEB is the SOTA on both CUB and NABirds.
5.5 Finetuning the pre-trained PEEB on CUB-200 yields a competitive explainable classifier in supervised learning
After finding that PEEB performs well in both GZSL (Sec. 5.3) and ZSL settings (Sec. 5.4), here we test finetuning the pre-trained PEEB on CUB-200. That is, we compare PEEB against SOTA explainable classifiers in the supervised learning setting to gain insights into our method’s adaptability to downstream tasks.
Experiment To understand the impact of pre-training and image resolution, we test finetuning three different PEEB variants: (1) PEEB initialized from OWL-ViT without pre-training on Bird-11K; (2) PEEB initialized from OWL-ViT with pre-training (described in Sec. 5.2); and (3) PEEB initialized from OWL-ViT with pre-training. We take each PEEB model and finetune all components on CUB-200, for 30 epochs with a batch size of 30, a learning rate of . Detailed hyperparameters are in Table A2.
Results Without pre-training, PEEB reaches 77.80% top-1 accuracy on CUB-200. Yet, first pre-training on Bird-11K and then finetuning on CUB yields 86.73%, the best among all explainable classifiers (Table 5b–c). Besides, pre-training PEEB from the higher-resolution OWL-ViT results in a further gain of +2.07 (86.73% 88.80%), which is intuitive since fine-grained classification is known to benefit from higher resolutions.
For a complete assessment, we compare and find PEEB to underperform SOTA standard, black-box classifiers by a few points (Table 5a).
| Methods | Model size | Backbone | Acc (%) |
| (a) SOTA black-box classifiers | |||
@@
| ViT-Net (2022a) | 26M | DeiT-S | 90.10 |
| (b) Concept-bottleneck classifiers | |||
| CBM Koh et al. (2020) | 11M | ResNet-18 | 80.10 |
| CPM Panousis et al. (2023) | 155M | ViT-B/16 | 72.00 |
| CDM Oikarinen et al. (2023) | 155M | ViT-B/16 | 74.31 |
| LaBo Yang et al. (2023) | 427M | ViT-L/14 | 81.90 |
| (c) Part-based, explainable classifiers | |||
@@
| ProtoTree (2021) | 92M | ResNet-50 | 82.20 |
| TesNet (2021) | 79M | DenseNet-121 | 84.80 |
@@
| PEEB (ours) | 155M | ||
| pre-training + finetuning only | 155M | OWL-ViT | 77.80 |
| pre-training + finetuning | 155M | OWL-ViT | 86.73 |
| pre-training + finetuning | 155M | OWL-ViT | 88.80 |
5.6 Applying PEEB to
dog identification
We have found that our pre-training dataset construction and PEEB performs well for bird identification. By design, our method is not specific to birds but is instead applicable to any fine-grained classification domains assuming that the object is decomposable into parts. Here, we show that our method performs well on dog image classification as well.
Pre-training dataset construction First, we define a set of six dog parts that humans use to identify dog species. We use all 4 dog parts defined by PartImageNet He et al. (2022b)—head, body, legs, and tail—and two more parts—muzzle and ears—based on our manual image examination.
We combine ImageNet-21K and Stanford Dogs-120 into Dog-140, our large-scale pre-training dataset spanning 140 dog species (details in Sec. D.2). For each class, we prompt GPT-4 to get the descriptors for 6 parts. For each image in Dog-140, we run OWL-ViT to detect the corresponding boxes for 6 pre-defined parts.
Experiment Following the supervised learning experiment in Sec. 5.5, we first we pre-train PEEB (initialized from OWL-ViT) on Dog-140 and then further finetune it on Dogs-120.
Results Finetuning PEEB on Dogs-120 from OWL-ViT without pre-training on Dog-140 results in a 74.17% top-1 accuracy on Dogs-120 (Table 6b). In contrast, pre-training on Dog-140 only without finetuning results in much better Dogs-120 accuracy of 87.38%. That is, our contrastive pre-training helps model generalize (in a GZSL setting) while directly finetuning on Dogs-120 perhaps yields an overfitting model. Yet, pre-training and then finetuning reaches the best supervised learning accuracy of 92.20%, which is SOTA among all explainable models reported on Dogs-120.
Besides, PEEB offers novel, editable image-text grounding explanations (see Fig. 5).
| Methods | Model size | Backbone | Acc (%) |
| (a) SOTA black-box classifiers | |||
| TransFG (2022a) | 86M | ViT-B/16 | 92.30 |
| ViT-Net (2022b) | 86M | DeiT-B | 93.60 |
| SR-GNN (2022) | 32M | Xception | 97.00 |
| (b) Explainable methods | |||